OpenStudy (anonymous):

Solve the equation -3 cos t=1 in the interval from 0 < theta < 2pi. Round to the nearest hundredth.

OpenStudy (campbell_st):

so you need to divide both sides of the equation by -3 which results in \[\cos(t) = \frac{-1}{3}\] so you need to know there are 2 solutions, 2 quadrants where t is negative so just find \[\cos(t) = \frac{1}{3}\] so just solve for t as a 1st quadrant angle then use 2nd quadrant 180 - t 3rd quadrant 180 + t to find the 2 solutions